10. For H2(g) + CO2(g) ←→CO(g) + H2O(g), AH 34.78 kJ/mol at 1000 K. If Kp = 0.236 at 800 K, calculate Kp at 1200 K, assuming that 4H does not vary significantly with temperature.
Answer #1

Q10

Apply van Hott

ln(Kp2/Kp1) = -H/R*(1/T1-1/T2)

ln(Kp2/0.236) = 347800/8.314*(1/800-1/1000)

ln(Kp2/0.236) = 10.458

Kp = 0.236*exp(10.458)

KP =8217.9 at 1000K

which makes sense since HRxn is endothermic, meaning that increase in T should increasE Kp